Contents

1 A section 1

1 A section

Here we will put some text, just for a test, nothing peculiar.

1. This will appear as item number one;

2. This will be item number two; the item command tells L^{A}TEX that we are starting a different point of the list; L^{A}TEX automatically inserts the correct number;

3. First-level lists use numbers when enumerating, bullets when listing;

• This is a nested list, bulleted; if it was a nested enumeration, you would get letters;

• Let’s now end this itemization;

4. Ok, let’s end this enumeration too;

The previous blank line tells TEX to end the previous paragraph and begin a new one. TEX then add necessary space between paragraphs, and indents the first line of each paragraph. Note that the linebreaks within each paragraph do not correspond to the linebreaks in the source: TEX reads material until he completes a paragraph (empty line), then it typesets the whole paragraph with optimum inter-word space, hyphenating where necessary (it will use english hyphenation rules thanks to the use of the Babel package).

We will now put here some math, like x = 0 or a < 1 < b; spaces in maths are not important, not even in so-called “displayed” math:

f(x) = \frac{sin x + cos x − \frac{1−cos x}{sin x}}{2x + 3x^{2}}

TEX will decide the correct spacing depending on the meaning of the tokens (if they are variables or relations or funcions etc). You can force the meaning, if you want, so you still keep full control on what is typeset and how.